Feather Friends, Inc., distributes a high-quality wooden birdhouse that sells for $80 per unit. Variable expenses are $40.00 per unit, and fixed expenses total $180,000 per year. Its operating results for last year were as follows:

Sales $ 2,080,000
Variable expenses 1,040,000
Contribution margin 1,040,000
Fixed expenses 180,000
Net operating income $ 860,000

Required:

Answer each question independently based on the original data:

1. What is the product's CM ratio?

2. Use the CM ratio to determine the break-even point in dollar sales.

3. Assume this year’s unit sales and total sales increase by 56,000 units and $4,480,000, respectively. If the fixed expenses do not change, how much will net operating income increase?

4-a. What is the degree of operating leverage based on last year's sales?

4-b. Assume the president expects this year's unit sales to increase by 15%. Using the degree of operating leverage from last year, what percentage increase in net operating income will the company realize this year?

5. The sales manager is convinced that a 13% reduction in the selling price, combined with a $70,000 increase in advertising, would increase this year's unit sales by 25%.

a. If the sales manager is right, what would be this year's net operating income if his ideas are implemented?

b. If the sales manager's ideas are implemented, how much will net operating income increase or decrease over last year?

6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $2.00 per unit. He thinks that this move, combined with some increase in advertising, would increase this year's unit sales by 25%. How much could the president increase this year's advertising expense and still earn the same $860,000 net operating income as last year?

Complete this question by entering your answers in the tabs below.

  • Req 1
  • Req 2
  • Req 3
  • Req 4A
  • Req 4B
  • Req 5A
  • Req 5B
  • Req 6

What is the product’s CM ratio?

CM ratio %

Use the CM ratio to determine the break-even point in dollar sales. (Do not round intermediate calculations.)

Break-even point in dollar sales

Use the CM ratio to determine the break-even point in dollar sales. (Do not round intermediate calculations.)

Break-even point in dollar sales

What is the degree of operating leverage based on last year's sales? (Round intermediate calculations and final answers to 2 decimal places.)

Degree of operating leverage

Assume the president expects this year's unit sales to increase by 15%. Using the degree of operating leverage from last year, what percentage increase in net operating income will the company realize this year? (Round intermediate calculations and final answer to 2 decimal places.)

Net operating income increases by %

The sales manager is convinced that a 13% reduction in the selling price, combined with a $70,000 increase in advertising, would increase this year's unit sales by 25%. If the sales manager is right, what would be this year's net operating income if his ideas are implemented? (Do not round intermediate calculations.)

Net operating income (loss)

The sales manager is convinced that a 13% reduction in the selling price, combined with a $70,000 increase in advertising, would increase this year's unit sales by 25%. If the sales manager's ideas are implemented, how much will net operating income increase or decrease over last year? (Negative amounts should be input with a minus sign.)

Increase (decrease) to net operating income

The president does not want to change the selling price. Instead, he wants to increase the sales commission by $2.00 per unit. He thinks that this move, combined with some increase in advertising, would increase this year's unit sales by 25%. How much could the president increase this year's advertising expense and still earn the same $860,000 net operating income as last year? (Do not round intermediate calculations.)

Show less

The amount by which advertising can be increased is
Answer #1

Contribution Margin Ratio = Contribution Margin/Sales

= 1,040,000/2,080,000

= 50%

Break even point in sales dollars = Fixed costs/Contribution Margin Ratio

= 180,000/50%

= $360,000

Increase in operating income = increase in contribution margin since fixed costs remain same

= 4,480,000*50%

= $2,240,000

4-a Degree of Operating leverage = Contribution Margin/Net Operating income

= 1,040,000/860,000

= 1.2093

4-b % increase in Income = % increase in sales*Degree of operating leverage

= 15%*1.2093 = 18.1395%

i.e. 18.14%

5a. Net Operating income = (69.6 – 40)*26000*1.25 – 180,000-70,000

= $712,000

b.Decrease in Net income = -$148,000

6.Desired Operating Income = $860,000

Add: Fixed costs other than advertising= 180,000

Total = $1,040,000

Contribution Margin = (80-40-2)*26,000*1.25 = $1,235,000

Increase in advertising expense = 1,235,000-1,040,000 = $195,000